OpenStudy (anonymous):

The pyramid shown has a square base with sides (3y-4)cm and a vertical height of 6 cm.The total volume of the pyramid is 32cm^3. @sammixboo

OpenStudy (anonymous):

i got 18y^2-48y=0

OpenStudy (anonymous):

this is my working

sammixboo (sammixboo):

Wait what are you trying to find?

OpenStudy (anonymous):

write down a quadratic equation in terms of y. sorry

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }(6)(3y-4)^2=32\]

OpenStudy (anonymous):

\[2\left[ (3y-4)(3y-4) \right]=32\]

sammixboo (sammixboo):

I believe that is correct from what I did. I just want a second opinion

OpenStudy (anonymous):

\[2(9y^2-24y+16)=32\] \[18y^2-48y+32=32\]

sammixboo (sammixboo):

Yes

sammixboo (sammixboo):

I believe you are correct

OpenStudy (anonymous):

thnx @sammixboo :)

sammixboo (sammixboo):

No problem :)